If the ratio of the crcles A and B is 4. WHat is the ratio of the radii of A and B?
Answer by stanbon(75887)

If the ratio of the circles A and B is 4.
What is the ratio of the radii of A and B?
-----
Let the radius of A be "x".
Let the radius of B be "y".
-----
Equation:
(pi)a^2/(pi/b^2) = 4/1
a^2/b^2 = 4/1
(a/b) = 2/1
-----------------
Cheers,
Stan H.